Find the slope of the curve and the equation of tangent line of the parametric equation to the given point.



5. x=√t , y = 2t + 4 , t=1



1
Expert's answer
2021-11-29T14:13:37-0500

Solution;

To find the equation of the tangent line, we need the slope "m=\frac{dy}{dx}" and the point of tangency "(x_o,y_o)"

Then the equation is the usual;

"(y-y_o)=m(x-x_o)"

"x=\sqrt t,y=2t+4,t=1"

So we compute;

"\frac{dx}{dt}=\frac{1}{\sqrt t}"

"\frac{dy}{dt}=2"

Therefore;

"\frac{dy}{dx}=\frac{dy}{dt}.\frac{dt}{dx}=2.\sqrt t=2\sqrt t"

Now put t=1;

"m=\frac{dy}{dx}=2\sqrt 1=2"

Also ,at t=1, the original equations give;

"x_o=\sqrt t=\sqrt 1=1"

"y_o=2t+4=2(1)+4=6"

Now we put in the info for the tangent line:

"y-y_o=m(x-x_o)"

"y-6=2(x-1)"

"y=2x+4"